# Weak expansion properties and a large deviation principle for coarse expanding conformal systems

### v2

@@ 5.3. Proof of Theorems A and B
We first verify the last statement. By [Expans], there exists $N\in\mathbb{N}$ such that for each $V^{\prime}\in\mathcal{U}_{N}$ there exists $V\in\mathcal{U}_{0}$ such that $V^{\prime}\subseteq V$. Then for each $k\in\mathbb{N}_{0}$ and each $U^{\prime}\in\mathcal{U}_{(k+1)N}$, there exists $U_{0}\in\mathcal{U}_{0}$ such that $f^{kN}(U^{\prime})\subseteq U_{0}$. Since $U^{\prime}$ is connected, there exists $U\in\mathcal{U}_{kN}$ such that $U$ is a connected component of $f^{-kN}(U_{0})$ and $U^{\prime}\subseteq U$. Thus $\mathcal{U}_{(k+1)N}$ is a refinement of $\mathcal{U}_{kN}$. Next, we check the Assumptions in Section 4 for $f^{n}|_{\mathfrak{X}_{n}}\colon(\mathfrak{X}_{n},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{kn}\}_{k\in\mathbb{N}_{0}}$ one by one. (1) $\mathfrak{X}_{n}$ is Hausdorff since $\mathfrak{X}_{0}$ is Hausdorff. For each $x\in\mathfrak{X}_{n}$, choose a compact neighborhood $U_{x}$ of $x$ in $\mathfrak{X}_{0}$. Since $f$ is proper, $f^{-n}(U_{x})$ is a compact neighborhood of $x$ in $\mathfrak{X}_{n}$, so $\mathfrak{X}_{n}$ is locally compact. Since $\mathfrak{X}_{n}=f^{-n}(\mathfrak{X}_{0})$ is open, for each $x\in\mathfrak{X}_{n}$ and each neighborhood $U$ of $x$ in $\mathfrak{X}_{n}$, $U$ is an open subset of $\mathfrak{X}_{0}$. By the local connectivity of $\mathfrak{X}_{0}$, there exists a connected neighborhood $V$ of $x$ such that $V\subseteq U\subseteq\mathfrak{X}_{n}$, thus $\mathfrak{X}_{n}$ is locally connected. Let $\mathfrak{C}_{n}$ be an arbitrary connected component in $\mathfrak{X}_{n}$ and $\mathfrak{C}_{0}$ be the connected component in $\mathfrak{X}_{0}$ which contains $f^{n}(\mathfrak{C}_{n})$. Then $\mathfrak{C}_{n}$ is a connected component of $f^{-n}(\mathfrak{C}_{0})$. By Proposition 3.5, $f^{n}|_{\mathfrak{X}_{n}}\colon{\mathfrak{X}_{n}\rightarrow\mathfrak{X}_{0}}$ is finite-to-one, open, closed and surjective. Since $\mathfrak{X}_{0}$ is locally connected, $\mathfrak{C}_{0}$ is open. Then by Proposition 3.6, ${f^{n}|_{\mathfrak{C}_{n}}\colon\mathfrak{C}_{n}\rightarrow\mathfrak{C}_{0}}$ is surjective. Since $\mathfrak{X}_{0}$ has finitely many connected components and $\deg(f^{n})<+\infty$, we get that $\mathfrak{X}_{n}$ has finitely many connected components. (2) Since $f$ is continuous, $\mathfrak{X}_{n}=f^{-(n-1)}(\mathfrak{X}_{1})$ is open. Since $\overline{\mathfrak{X}_{n}}$ is compact in $\mathfrak{X}_{n-1}$, $\overline{\mathfrak{X}_{n}}$ is compact in $\mathfrak{X}_{0}$. (3) Let $\deg(f)=d>1$. The composition of two FBC maps is an FBC map, and the degree multiplies, so $\deg(f^{n})=d^{n}$. Since for each $k\in\mathbb{N}$, $f|_{\mathfrak{X}_{k}}\colon\mathfrak{X}_{k}\rightarrow\mathfrak{X}_{k-1}$ is an FBC map of degree $d$, $f^{n}|_{\mathfrak{X}_{n}}$ is an FBC map of degree $d^{n}$. The repellor of $f^{n}|_{\mathfrak{X}_{n}}\colon\mathfrak{X}_{n}\rightarrow\mathfrak{X}_{0}$ is $X=\bigcap_{k\in\mathbb{N}}\overline{\mathfrak{X}_{k}}=\bigcap_{k\in\mathbb{N}}\overline{\mathfrak{X}_{kn}}$. Since $f|_{X}\colon X\rightarrow X$ is an FBC map of degree $d$, $f^{n}|_{X}=(f|_{X})^{n}$ is an FBC map of degree $d^{n}$. (4) By definition $\mathcal{U}_{n}$ is a finite cover of $X$ by connected open subsets of $\mathfrak{X}_{n}$ and for each $k\in\mathbb{N}_{0}$ the elements of $\mathcal{U}_{(k+1)n}$ are preimages under $f^{n}$ of elements in $\mathcal{U}_{kn}$. Axioms [Expans], [Irred], [Deg], [Round], and [Diam] hold for the sequence $\{\mathcal{U}_{kn}\}_{k\in\mathbb{N}}$, which is a straightforward consequence of the definitions. ∎
@@
Next, we prove that a metric CXC system is not $h$-expansive if there exists at least one branch point in the repellor.
@@
Proof of Theorem B.
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4, and $x_{0}\in X\cap B_{f}$. Denote $f_{X}\coloneqq f|_{X}$.
@@
    By [Deg], for each $x\in X$, the forward orbit $\{f^{n}(x)\}_{n\in\mathbb{N}_{0}}$ contains at most $\lceil\log_{2}p\rceil$ many branch points. So we can assume that for each $n\in\mathbb{N}$ and each $x\in f^{-n}(x_{0})$, $\deg(f;x)=1$.
Denote $\mathcal{W}_{n}\coloneqq\{U\cap X:U\in\mathcal{U}_{n}\}$ for each $n\in\mathbb{N}_{0}$. By Lemma 5.11, $\{\mathcal{W}_{n}\}_{n\in\mathbb{N}_{0}}$ is a refining sequence of open covers of $X$. Thus it follows from (5.1) and (5.2) in Definition 5.2 that it suffices to prove that
@@
Let $k\coloneqq\deg(f;x_{0})$. Then $k>1$. By [Expans], for sufficiently large $m\in\mathbb{N}$, there exists $U_{0}^{m}\in\mathcal{U}_{m}$ such that $x_{0}\in U_{0}^{m}$ and that $f|_{U_{0}^{m}}\colon U_{0}^{m}\rightarrow U_{0}^{m-1}$ is an FBC map of degree $k$. Fix such $m$ and $U_{0}^{m}$, and denote $U_{0}^{m-1}\coloneqq f(U_{0}^{m})$.
@@
[…]
(5.19) $\delta\coloneqq\min\bigl{\{}d\bigl{(}\widetilde{U}_{i},\widetilde{U}_{j}\bigr{)}:1\leq i<j\leq k\bigr{\}}>0.$
Fix an arbitrary number
@@
[…]
@@
[…]
Case 1. Assume $0\leq r<t$.
By our induction hypothesis, $A_{(t+1)s+r}\in V^{m+r}$. For each $x\in A_{n}$, choose one point $x^{\prime}\in V^{m+r+1}$ such that $f(x^{\prime})=x$. Define $A_{(t+1)s+r+1}$ to be the collections of all such $x^{\prime}$ corresponds to $x\in A_{(t+1)s+r}$.
@@
[…]
$A_{(t+1)s+r+1}\subseteq\widehat{U}^{m}\cap\bigcap_{i=1}^{(t+1)s+r+1}f^{-i}(\widehat{U}_{i-1}^{m})\in\bigvee_{i=0}^{(t+1)s+r+1}f^{-i}(\mathcal{U}_{m}).$
@@
Next, we verify condition (3). By definition, $\operatorname{card}(A_{(t+1)s+r+1})=\operatorname{card}(A_{(t+1)s+r})$. When $0\leq r<t$, we have $\bigl{\lfloor}\frac{(t+1)s+r+1}{t+1}\bigr{\rfloor}=s=\bigl{\lfloor}\frac{(t+1)s+r}{t+1}\bigr{\rfloor}$, and condition (3) is verified.
[…]
$\{f(x),\,f(y)\}\subseteq\bigcap_{i=0}^{(t+1)s+r}f^{-i}(\widehat{U}_{i+1}^{l})\in\bigvee_{i=0}^{(t+1)s+r}f^{-i}(\mathcal{U}_{l}),$
@@
[…]
@@
[…]
@@
[…]
@@
So for such $A\in\bigvee_{i=0}^{n}f_{X}^{-i}(\mathcal{W}_{m})$ that contains $A_{n}$, if $\mathcal{B}\subseteq\bigvee_{i=0}^{n}f_{X}^{-i}(\mathcal{W}_{l})$ satisfies $A\subseteq\bigcup\mathcal{B}$, then $\operatorname{card}(\mathcal{B})\geq\operatorname{card}(A_{n})=k^{\lfloor\frac{n}{t+1}\rfloor}$.
@@
(5.20) $H\biggl{(}\bigvee_{i=0}^{n-1}f_{X}^{-i}(\mathcal{W}_{l})\,\bigg{|}\,\bigvee_{j=0}^{n-1}f_{X}^{-j}(\mathcal{W}_{m})\biggr{)}\geq\log k^{\lfloor\frac{n-1}{t_{m}+1}\rfloor}.$
@@
$\geq$ $\liminf_{l\to+\infty}\liminf_{n\to+\infty}\frac{1}{n}\log k^{\lfloor\frac{n-1}{t_{m}+1}\rfloor}=\frac{\log k}{t_{m}+1}>0.$
Therefore, the map $f|_{X}$ is not $h$-expansive. ∎

### v3

@@ 5.3. Proofs of Theorems A and B
We first verify the last statement. By [Expans], there exists $N\in\mathbb{N}$ such that for each $V^{\prime}\in\mathcal{U}_{N}$ there exists $V\in\mathcal{U}_{0}$ such that $V^{\prime}\subseteq V$. Then for each $k\in\mathbb{N}_{0}$ and each $U^{\prime}\in\mathcal{U}_{(k+1)N}$, there exists $U_{0}\in\mathcal{U}_{0}$ such that $f^{kN}(U^{\prime})\subseteq U_{0}$. Since $U^{\prime}$ is connected, there exists $U\in\mathcal{U}_{kN}$ such that $U$ is a connected component of $f^{-kN}(U_{0})$ and $U^{\prime}\subseteq U$. Thus $\mathcal{U}_{(k+1)N}$ is a refinement of $\mathcal{U}_{kN}$.
Next, we check the Assumptions in Section 4 for $f^{n}|_{\mathfrak{X}_{n}}\colon(\mathfrak{X}_{n},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{kn}\}_{k\in\mathbb{N}_{0}}$ one by one.
(1) $\mathfrak{X}_{n}$ is Hausdorff since $\mathfrak{X}_{0}$ is Hausdorff.
For each $x\in\mathfrak{X}_{n}$, choose a compact neighborhood $U_{x}$ of $f^{n}(x)$ in $\mathfrak{X}_{0}$. Since $f$ is proper, $f^{-n}(U_{x})$ is a compact neighborhood of $x$ in $\mathfrak{X}_{n}$, so $\mathfrak{X}_{n}$ is locally compact.
Since $\mathfrak{X}_{n}=f^{-n}(\mathfrak{X}_{0})$ is open, for each $x\in\mathfrak{X}_{n}$ and each neighborhood $U$ of $x$ in $\mathfrak{X}_{n}$, $U$ is an open subset of $\mathfrak{X}_{0}$. By the local connectivity of $\mathfrak{X}_{0}$, there exists a connected neighborhood $V$ of $x$ such that $V\subseteq U\subseteq\mathfrak{X}_{n}$, thus $\mathfrak{X}_{n}$ is locally connected.
Let $\mathfrak{C}_{n}$ be an arbitrary connected component in $\mathfrak{X}_{n}$ and $\mathfrak{C}_{0}$ be the connected component in $\mathfrak{X}_{0}$ which contains $f^{n}(\mathfrak{C}_{n})$. Since $\mathfrak{X}_{0}$ is locally connected, $\mathfrak{C}_{0}$ is open. Then $\mathfrak{C}_{n}$ is a connected component of $f^{-n}(\mathfrak{C}_{0})$. By Proposition 3.5, $f^{n}|_{\mathfrak{X}_{n}}\colon{\mathfrak{X}_{n}\rightarrow\mathfrak{X}_{0}}$ is finite-to-one, open, closed, and surjective. Then by Proposition 3.6, $f^{n}|_{\mathfrak{C}_{n}}\colon\mathfrak{C}_{n}\rightarrow\mathfrak{C}_{0}$} is surjective. Since $\mathfrak{X}_{0}$ has finitely many connected components and $\deg(f^{n})<+\infty$, we get that $\mathfrak{X}_{n}$ has finitely many connected components.
(2) Since $f$ is continuous, $\mathfrak{X}_{n}=f^{-(n-1)}(\mathfrak{X}_{1})$ is open. Since $\overline{\mathfrak{X}_{n}}$ is compact in $\mathfrak{X}_{n-1}$, $\overline{\mathfrak{X}_{n}}$ is compact in $\mathfrak{X}_{0}$.
(3) Let $\deg(f)=d>1$. The composition of two FBC maps is an FBC map, and the degree multiplies, so $\deg(f^{n})=d^{n}$.
Since for each $k\in\mathbb{N}$, $f|_{\mathfrak{X}_{k}}\colon\mathfrak{X}_{k}\rightarrow\mathfrak{X}_{k-1}$ is an FBC map of degree $d$, $f^{n}|_{\mathfrak{X}_{n}}$ is an FBC map of degree $d^{n}$. The repellor of $f^{n}|_{\mathfrak{X}_{n}}\colon\mathfrak{X}_{n}\rightarrow\mathfrak{X}_{0}$ is $X=\bigcap_{k\in\mathbb{N}}\overline{\mathfrak{X}_{k}}=\bigcap_{k\in\mathbb{N}}\overline{\mathfrak{X}_{kn}}$. Since $f|_{X}\colon X\rightarrow X$ is an FBC map of degree $d$, $f^{n}|_{X}=(f|_{X})^{n}$ is an FBC map of degree $d^{n}$.
(4) By definition $\mathcal{U}_{n}$ is a finite cover of $X$ by connected open subsets of $\mathfrak{X}_{n}$ and for each $k\in\mathbb{N}_{0}$ the elements of $\mathcal{U}_{(k+1)n}$ are preimages under $f^{n}$ of elements in $\mathcal{U}_{kn}$. Axioms [Expans], [Irred], [Deg], [Round], and [Diam] hold for the sequence $\{\mathcal{U}_{kn}\}_{k\in\mathbb{N}}$, which is a straightforward consequence of the definitions. ∎
@@
In what follows, we show that a metric CXC system is forward expansive if there is no branch point in the repellor. The proof is based on the observation that the distance between two points $x,\,y\in X$ can be estimated by the diameter of some good open set containing one of x,yx,\,y, and such estimation works well under iterations when $d(x,y)$ is small.
Theorem 5.16.
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. Assume that $X\cap B_{f}=\emptyset$. Then $f|_{X}$ is forward expansive.
Proof.
For each $x\in X$, since $\deg(f;x)=1$, there exists an open neighborhood $U_{x}$ of $x$, such that $f|_{U_{x}}$ is injective. Let $\mathcal{V}\coloneqq\{U_{x}:x\in X\}$. By Lemma 5.9, we may fix $n_{0}\in\mathbb{N}$, such that for all $n,\,k\in\mathbb{N}_{0}$ with $n\geqslant n_{0}$ and each $U^{n}\in\mathcal{U}_{n}$, there exists $V\in\mathcal{V}$ for which $\operatorname{B}_{n}^{k}(U^{n})\subseteq V$. Thus, $f^{n-n_{0}}|_{\operatorname{B}_{n}^{k}(U^{n})}$ is injective for all $n,\,k\in\mathbb{N}_{0}$ with $n\geqslant n_{0}$ and each $U^{n}\in\mathcal{U}_{n}$.
Consider
$\delta\coloneqq\min\biggl{\{}\frac{C^{\prime\prime}\lambda^{n_{0}}}{2K},\,\frac{C^{\prime\prime}\lambda^{n_{0}}}{2\rho_{+}(K)}\biggr{\}},$
where $C^{\prime\prime}>0$, $\lambda\in(0,1)$ are the constants in Lemma 5.7, $K>1$ is the constant in Proposition 3.10, and $\rho_{+}$ is the increasing embedding in [Round].
Fix arbitrary $\in X, xyx\neq y$. Suppose that $d(x,y)<\delta\leqslant\frac{C^{\prime\prime}\lambda^{n_{0}}}{2K}$. Then by Proposition 3.10 (iv), there exists $U^{n_{0}}\in\mathcal{U}_{n_{0}}$ containing x,yx,\,y. Put
$n\coloneqq\max\{l\in\mathbb{N}:\mbox{there exists }U\in\mathcal{U}_{l}\mbox{ such that }x,\,y\in U\},$
then we have $n\geqslant n_{0}$.
Suppose that $U\in\mathcal{U}_{n}$ and $x,\,y\in U$. By Proposition 3.10 (ii), there exists $U^{\prime}\in\mathcal{U}_{n+1}$ such that $x\in U$, and that $\operatorname{Round}(U^{\prime},x)<K$. Here $K$ is the constant in Proposition 3.10. By the choice of $n$, we have that $y\notin U^{\prime}$.
Denote $k\coloneqq n-n_{0}$. Since $U\cup U^{\prime}\subseteq\operatorname{B}_{n}^{1}(U)$, the map $f^{k}|_{U\cap U^{\prime}}$ is injective. Thus, $f^{k}(y)\notin f^{k}(U^{\prime})$.
By [Round], we have
$\operatorname{Round}\bigl{(}f^{k}(U^{\prime}),f^{k}(x)\bigr{)}<\rho_{+}(\operatorname{Round}(U^{\prime},x))<\rho_{+}(K).$
Denote $L\coloneqq\operatorname{Round}\bigl{(}f^{l}(U^{\prime}),f^{k}(x)\bigr{)}$. Then, there exists $r>0$, such that
$B\bigl{(}f^{k}(x),r\bigr{)}\subseteq f^{k}(U^{\prime})\subseteq\overline{B\bigl{(}f^{k}(x),Lr\bigr{)}}.$
It is clear that
$r\geqslant\frac{\operatorname{diam}(f^{k}(U^{\prime}))}{2L}\geqslant\frac{\operatorname{diam}(f^{k}(U^{\prime}))}{2\rho_{+}(K)}.$
Since $f^{k}(y)\notin f^{k}(U^{\prime})$, we have that
$d\bigl{(}f^{k}(x),f^{k}(y)\bigr{)}\geqslant r\geqslant\frac{\operatorname{diam}(f^{k}(U^{\prime}))}{2\rho_{+}(K)}\geqslant\frac{C^{\prime\prime}\lambda^{n_{0}}}{2\rho_{+}(K)}\geqslant\delta.$
Since $x,\,y\in X$ are arbitrary, $f|_{X}$ is forward expansive. ∎
Remark.
Note that $f|_{X}$ cannot be forward expansive if there exists a branch point in $X$.
Next, we prove that a metric CXC system is $h$-expansive if and only if there is no branch point in the repellor. Here is the main idea of the proof. Given a sufficiently large $m\in\mathbb{N}$, we construct recursively a sequence $\{A_{n}\}_{n\in\mathbb{N}}$ of subsets of $X$ such that each $A_{n}$ is contained in an element in $\bigvee_{i=0}^{n}f^{-i}(\mathcal{W}_{m})$. We prove by induction that for each sufficiently large $l\in\mathbb{N}$, the number of elements in $\bigvee_{j=0}^{n}f^{-j}(\mathcal{W}_{l})$ needed to cover $A_{n}$ is bounded from below, and such lower bound leads to the conclusion that $h(f|\mathcal{W}_{m})>0$.
@@
Theorem 5.18.
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4. Suppose that $B_{f}\cap X\neq\emptyset$. Then $f|_{X}$ is not $h$-expansive.
Proof.
Consider $x_{0}\in B_{f}\cap X$.
@@
    By [Deg], for each $x\in X$, the forward orbit $\{f^{n}(x)\}_{n\in\mathbb{N}_{0}}$ contains at most $\lceil\log_{2}p\rceil$ many branch points. Thus, there exists a branch point whose preimages under iterations of $f$ contain no branch point. We may assume that $\deg(f;x)=1$ for each $n\in\mathbb{N}$ and each $x\in f^{-n}(x_{0})$.
By Lemma 5.13, $\{\mathcal{W}_{n}\}_{n\in\mathbb{N}_{0}}$ is a refining sequence of open covers of $X$. Thus, it follows from (5.1) and (5.3) in Definition 5.2 that it suffices to prove that
@@
Let $k\coloneqq\deg(f;x_{0})$. Then $k>1$. By [Expans], for each sufficiently large $m\in\mathbb{N}$, there exists $U_{0}^{m}\in\mathcal{U}_{m}$ such that $x_{0}\in U_{0}^{m}$ and that $f|_{U_{0}^{m}}\colon U_{0}^{m}\rightarrow f(U_{0}^{m})$ is an FBC map of degree $k$. Fix such $m$ and $U_{0}^{m}$, and denote $U_{0}^{m-1}\coloneqq f(U_{0}^{m})$.
@@
    For all $U,\,U^{\prime}\in\mathcal{U}_{m^{\prime}+1}$, if $f(U)=f(U^{\prime})=U_{0}^{m^{\prime}}$ and $z_{i}\in U$, $z_{j}\in U^{\prime}$ for some $i,\,j\in\{1,\,2,\,\cdots,\,k\}$ with $i\neq j$, then $d(U,U^{\prime})\coloneqq\inf\{d(x,y):x\in U,\,y\in U^{\prime}\}>0$.
For each $i\in\{1,\,2,\,\cdots,\,k\}$, choose $\widetilde{U}_{i}\in\mathcal{U}_{m^{\prime}+1}$ such that $f\bigl{(}\widetilde{U}_{i}\bigr{)}=U_{0}^{m^{\prime}}$ and that $z_{i}\in\widetilde{U}_{i}$. Then $d\bigl{(}\widetilde{U}_{i},\widetilde{U}_{j}\bigr{)}>0$ for all $1\leqslant i<j\leqslant k$. We denote
(5.20) $\delta\coloneqq\min\bigl{\{}d\bigl{(}\widetilde{U}_{i},\widetilde{U}_{j}\bigr{)}:1\leqslant i<j\leqslant k\bigr{\}}>0.$
Fix an arbitrary natural number
@@
By Proposition 3.8, there exists $t\in\mathbb{N}$ and $V^{m+t}\in\mathcal{U}_{m+t}$}, such that $f^{t}(V^{m+t})=U_{0}^{m}$, and that $V^{m+t}\subseteq U_{0}^{m^{\prime}}$. We denote $V^{m+r}\coloneqq f^{t-r}(V^{m+t})$ for each $r\in\{0,\,1,\,\dots,\,t-1\}$. In particular, $V^{m}=U_{0}^{m}$.
@@
Make the induction hypothesis that $A_{n}$ is defined and satisfies conditions (1) through (5) for each $n\in\{0,\,1,\,\dots,\,(t+1)s+r\}$, where $s\in\mathbb{N}_{0}$ and $r\in\{0,\,1,\,\dots,\,t\}$. We construct $A_{(t+1)s+r+1}$ in the following cases depending on $r$, and show that $A_{(t+1)s+r+1}$ satisfies conditions (1) through (5) (with $n\coloneqq(t+1)s+r+1$).
Case 1. Assume $0\leqslant r<t$.
By our induction hypothesis, $A_{(t+1)s+r}\in V^{m+r}$. For each $x\in A_{(t+1)s+r}$, choose one point $x^{\prime}\in V^{m+r+1}$ such that $f(x^{\prime})=x$. Define $A_{(t+1)s+r+1}$ to be the collection of all such $x^{\prime}$ that correspond to $x\in A_{(t+1)s+r}$.
@@
We first verify condition (1). By our induction hypothesis, $A_{(t+1)s+r}\subseteq A$ for some $A\in\bigvee_{i=0}^{(t+1)s+r}f^{-i}(\mathcal{U}_{m})$. Suppose that $A=\bigcap_{i=0}^{(t+1)s+r}f^{-i}\bigl{(}\widehat{U}_{i}^{m}\bigr{)}$ for some $\widehat{U}_{i}^{m}\in\mathcal{U}_{m}$ for each $i\in\{0,\,1,\,\dots,\,(t+1)s+r\}$. Since $A_{(t+1)s+r+1}\subseteq V^{m+r+1}$, and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ is a refining sequence, there exists $\widehat{U}^{m}\in\mathcal{U}_{m}$ such that $A_{(t+1)s+r+1}\subseteq V^{m+r+1}\subseteq\widehat{U}^{m}$. Thus
$A_{(t+1)s+r+1}\subseteq\widehat{U}^{m}\cap\bigcap_{i=1}^{(t+1)s+r+1}f^{-i}\bigl{(}\widehat{U}_{i-1}^{m}\bigr{)}\in\bigvee_{i=0}^{(t+1)s+r+1}f^{-i}(\mathcal{U}_{m}).$
@@
Next, we verify condition (3). By definition, $\operatorname{card}\bigl{(}A_{(t+1)s+r+1}\bigr{)}=\operatorname{card}\bigl{(}A_{(t+1)s+r}\bigr{)}$. When $0\leqslant r<t$, we have $\bigl{\lfloor}\frac{(t+1)s+r+1}{t+1}\bigr{\rfloor}=s=\bigl{\lfloor}\frac{(t+1)s+r}{t+1}\bigr{\rfloor}$, and condition (3) is verified.
Then, we verify condition (5). Suppose that there are $x,\,y\in A_{(t+1)s+r+1}$ with $x\neq y$, such that $\{x,\,y\}\subseteq B$ for some $B\in\bigvee_{i=0}^{(t+1)s+r+1}f^{-i}(\mathcal{U}_{l})$. Suppose $B=\bigcap_{i=0}^{(t+1)s+r+1}f^{-i}\bigl{(}\widehat{U}_{i}^{l}\bigr{)}$ for some $\widehat{U}_{i}^{l}\in\mathcal{U}_{l}$ for each $i\in\{0,\,1,\,\cdots,\,(t+1)s+r+1\}$. By definition $f(x),\,f(y)\in A_{(t+1)s+r}$ and $f(x)\neq f(y)$. However,
$\{f(x),\,f(y)\}\subseteq\bigcap_{i=0}^{(t+1)s+r}f^{-i}\bigl{(}\widehat{U}_{i+1}^{l}\bigr{)}\in\bigvee_{i=0}^{(t+1)s+r}f^{-i}(\mathcal{U}_{l}),$
@@
By our induction hypothesis $A_{(t+1)s+t}\subseteq V^{m+t}\subseteq U_{0}^{m^{\prime}}$. Fix an arbitrary number $i\in\{1,\,2,\,\dots,\,k\}$. For each $x\in A_{(t+1)s+t}$, choose one point $x^{\prime}\in\widetilde{U}_{i}$ such that $f(x^{\prime})=x$. Define $A^{i}_{(t+1)(s+1)}$ to be the collection of all such $x^{\prime}$ corresponding to some $x\in A_{(t+1)s+t}$.
@@
First, we verify condition (3). By construction, $\operatorname{card}\bigl{(}A^{i}_{(t+1)(s+1)}\bigr{)}=\operatorname{card}\bigl{(}A_{(t+1)s+t}\bigr{)}$. Since $A_{(t+1)(s+1)}^{i}\subseteq\widetilde{U}_{i}$ for each $i\in\{1,\,2,\,\dots,\,k\}$, and $\widetilde{U}_{i}\cap\widetilde{U}_{j}=\emptyset$ for each pair $i,\,j\in\{1,\,2,\,\dots,\,k\}$ with $i\neq j$, we have
@@
Next, we verify condition (5). Suppose that there are $x,\,y\in A_{(t+1)(s+1)}$ with $x\neq y$ such that $\{x,\,y\}\subseteq B$ for some $B\in\bigvee_{a=0}^{(t+1)(s+1)}f^{-a}(\mathcal{U}_{l})$. Suppose $B\subseteq U^{l}$ for some $U^{l}\in\mathcal{U}_{l}$. By Proposition 3.11 (i), $\operatorname{diam}(U^{l})\leqslant C^{\prime}d_{0}\theta^{l}$, so $d(x,\,y)<\delta$ since $l>\log_{\theta}\frac{\delta}{C^{\prime}d_{0}}$. Then by (5.20) there is some $i\in\{1,\,2,\,\dots,\,k\}$ such that $\{x,\,y\}\subseteq\widetilde{U}_{i}$, and by construction $f(x)\neq f(y)$. However, since $\{x,\,y\}\subseteq B$ and $B\in\bigvee_{a=0}^{(t+1)(s+1)}f^{-a}(\mathcal{U}_{l})$, we have $\{f(x),\,f(y)\}\subseteq S$ for some $S\in\bigvee_{a=0}^{(t+1)s+t}f^{-a}(\mathcal{U}_{l})$, which contradicts condition (5) for $A_{(t+1)s+t}$ in the induction hypothesis.
@@
So for such $A\in\bigvee_{i=0}^{n}f_{X}^{-i}(\mathcal{W}_{m})$ that contains $A_{n}$, if $\mathcal{B}\subseteq\bigvee_{i=0}^{n}f_{X}^{-i}(\mathcal{W}_{l})$ satisfies $A\subseteq\bigcup\mathcal{B}$, then $\operatorname{card}(\mathcal{B})\geqslant\operatorname{card}(A_{n})=k^{\lfloor\frac{n}{t+1}\rfloor}$.
@@
(5.21) $H\biggl{(}\bigvee_{i=0}^{n-1}f_{X}^{-i}(\mathcal{W}_{l})\,\bigg{|}\,\bigvee_{j=0}^{n-1}f_{X}^{-j}(\mathcal{W}_{m})\biggr{)}\geqslant\log k^{\lfloor\frac{n-1}{t_{m}+1}\rfloor}.$
@@
$\geqslant$ $\liminf_{l\to+\infty}\liminf_{n\to+\infty}\frac{1}{n}\log k^{\lfloor\frac{n-1}{t_{m}+1}\rfloor}=\frac{\log k}{t_{m}+1}>0.$
Therefore, the map $f|_{X}$ is not $h$-expansive, and the proof is complete. ∎
Here we give a quick proof of Theorem B.
Proof of Theorem B.
The proof follows from Remark 5.5, Theorems 5.16 and 5.18.
[…]